OpenStudy (anonymous):

The volume V of a sphere is four thirds π times the cube of its radius. Express the volume V of the sphere as a function of the radius r.

zepdrix (zepdrix):

Just translate the key words they give you. `is` means `equals`. Volume = 4/3(pi) times r^3. Yes? Simple as that. To show that Volume is a `function` of the radius, we can use function notation. Replace the V with V of r. \(\Large\rm V(r)=\dfrac{4}{3}\pi r^3\)

OpenStudy (anonymous):

Thanks!